OpenStudy (anonymous):

how to order fractions greatest to least 2/7 4/7 6/7

OpenStudy (anonymous):

hello

OpenStudy (anonymous):

study 23

OpenStudy (anonymous):

hello

OpenStudy (anonymous):

Think of arranging them on a number line. \(\ \huge \frac{7}{7} = 1 \). That means 6/7 is the largest because it is closest to one, or 7/7. Next, 4/7 is the second largest, and 2/7 is the least largest. So, the final order is \(\ \huge \frac{6}{7}, \frac{4}{7}, and \frac {2}{7} \).

OpenStudy (anonymous):

\(\ \huge Does that make sense? \)

OpenStudy (anonymous):

no its 2/7 4/7 6/7